What is the pH of a solution obtained after mixing 20 mL of 0.1 M HI solution with 10 mL of 0.2 M HCl?
Answer #1

both HI and HCl are strong acids. They ionize completely

since 1000ml= 1Lliters,

moles of HI = molarity* volume in liters= 0.1*20/1000 =0.002 and moles of HCl= 0.2*10/1000=0.002

HI +H2O--------->H3O++I-. moles of H3O+=0.002 and HCl+ H2O -------->H3O+ Cl-, moles of H3O= moles of HCl=0.002

total moles of H3O+=0.002+0.002= 0.004, volume of solution =20+10= 30ml, =30/1000Liters

[H3O+]= 0.004/(30/1000)= 0.133M

pH= -log[H3O+]= -log (0.133)= 0.875